Consider two jars, Jar M and Jar W. In Jar M, there are 3 balls numbered 0, 1, 2. In Jar W there are 3 balls numbered 1, 2, 3. A ball is drawn from Jar M, then a ball is drawn from Jar W. Define M as the number on the ball from Jar A and W the number on the ball drawn from Jar B. Let X = MW(M times W) and Y “ maxtW, Mu(Maximum of W and M). (a) Find the joint probability mass function of X and Y . (b) Are X and Y independent? Use the definition of independence to support your answer. (c) Compute the CorrpX, Y q (ie, Correlation coefficient of X and Y). (d) Find the probability mass function of X given Y “ 3 (ie, ppx|y “ 3q)
Answer #1

Soludon : Probability mass function of мs m0, 1, 2 Pin (m) { 0 otheowise is of w w 0,113 PMF Pw (W) 0 otherwise a x 1 x canP(x:2, Yzy ) - PIM 22, Wz1, Vey) ply=y) 2 it y 31, 2 13 if y - 3 o otherwise Joind PMF: for P(x=0, x=y) = P(=0, way) 1,2,3 92plx=4, Yay) - P(M=2,W=2, V=2) n y = 2 {: O Otherwise x 36, For PIXE6, Y=y) = P(M=2,W=3, Yey) p(x=6,x=y) 4:3 {: Ow Thus, joiDIX = 6 ): p(x26,Y=3 ) : 19 Plx:1) = 113 / 119 x-1,3,4,6 119 , 2 - 2 6 else Nolo are not in dependent P(x), Vry) # P (**») PI(ov XV). 19)/2) 44 - 38 9 - 2 3 • Co-A ( x + ), (ev (x, y) Var (x) Var (y) var (x) E lx?) - (F(x)) ? (1244+2*** *3*** *** *toa plx=7 14-3) - P(x, 1=3) PV =3) P(x=4 , Y-3) A13 Pix:+/y: 3 ) : 3 xta 220, 3,6 fore otherwise 113 ,01-0, 3,6 0 otherwise Ple